# Using High-Level Patterns to Estimate How Humans Predict a Robot will Behave

### v1

@@ Abstract
A human interacting with a robot often forms predictions of what the robot will do next. For instance, based on the recent behavior of an autonomous car, a nearby human driver might predict that the car is going to remain in the same lane. It is important for the robot to understand the human’s prediction for safe and seamless interaction: e.g., if the autonomous car knows the human thinks it is not merging — but the autonomous car actually intends to merge — then the car can adjust its behavior to prevent an accident. Prior works typically assume that humans make precise predictions of robot behavior. However, recent research on human-human prediction suggests the opposite: humans tend to approximate other agents by predicting their high-level behaviors. We apply this finding to develop a second-order theory of mind approach that enables robots to estimate how humans predict they will behave. To extract these high-level predictions directly from data, we embed the recent human and robot trajectories into a discrete latent space. Each element of this latent space captures a different type of behavior (e.g., merging in front of the human, remaining in the same lane) and decodes into a vector field across the state space that is consistent with the underlying behavior type. We hypothesize that our resulting high-level and course predictions of robot behavior will correspond to actual human predictions. We provide initial evidence in support of this hypothesis through a proof-of-concept user study.
@@ I Introduction
This concept of reasoning over another agent’s mental model is known as Theory of Mind (ToM) [5]. Specifically, in this paper we will focus on building a second-order ToM of the human (e.g., "the robot thinks about how human predicts the robot will act"). Existing works often focus only on a first-order ToM: modeling the human’s goal so they can provide better assistance [6], estimating their strategy in order to influence the human’s behavior [7, 8], or modeling their intent to improve collaboration between humans and robots [9, 10]. Prior research has also explored building second-order ToM for robots, for instance [11] models how robot’s actions affect a human’s behavior and leverage this to develop efficient and influential robot policy. When applied to our example, these works might expect the human to precisely predict each individual action of the autonomous car (see Figure 1, left). However, recent research on human-human interaction suggests that humans do not precisely anticipate the actions of other agents [12]; instead, humans appear to reason over more general and intuitive patterns of behavior. We therefore hypothesize that:
@@
In this paper, we apply our hypothesis to create an algorithmic, data-driven method to estimate how humans predict a robot will behave. Specifically, we develop an autoencoder that inputs recent human and robot behaviors and embeds these trajectories into a discrete latent space (Section IV). This learned latent space (a) inherently forces an information bottleneck that leads to coarse, human-like predictions, and (b) classifies the input trajectories into the discrete, high-level behaviors that were observed in the training dataset. We show, these discrete latent values often correspond to human-friendly patterns: e.g., merging in front of the human, remaining in the same lane, or merging behind the human (Section V). When we decode these discrete latent representations, the robot obtains a vector field of predictions over the environment’s state space. In practice, this enables the robot to estimate how the human thinks it will move right now (in its current state), and also during future timesteps (at other environment states). Our initial user study suggests that the second-order ToM predictions obtained using this vector field better align with actual human predictions than existing baselines (Section VI).
@@ II Related Work
Modeling Humans. Existing research in human-robot interaction has explored how robots can model human behavior. These works typically focus on estimating the human’s goal, objective, or intent. For example, in shared autonomy robots infer the human’s goals [13, 14], their intentions [6], even their skills [15] in order to assist humans. Alternatively, in multi-agent interaction robots form high-level representations of humans [7, 16, 17] to model their behavior (e.g., inferring whether human are aggressive or defensive drivers). Overall, these works form first-order models of the human (i.e., they estimate what the human is thinking). But they do not consider how the human reasons over the robot — i.e., how the human expects the robot to behave during interaction.
@@ III Problem Statement
Two Player Game. Formally, the human and robot are playing a two-player game. Let $s_{i}\in\mathcal{S}$ be the state of agent $i$, and let $a_{i}\in\mathcal{A}$ be the action of agent $i$. In what follows, we will use $i=1$ to refer to the robot and $i=2$ to refer to the human. Both agents (i.e., the human and robot) have policies that map the system state to actions. The robot’s policy $\pi_{1}(s)\rightarrow a_{1}$ determines the robot’s actions, and similarly the human’s policy $\pi_{2}(s)\rightarrow a_{2}$ determines the human’s actions. Within our problem setting the robot knows its own policy, but it does not know how the human is modeling the robot’s policy. On the other hand, the human is unaware of the robot’s policy $\pi_{1}$, though they have an understanding of their own predictions regarding the robot’s behavior.
@@ IV Extracting High-Level Behaviors for Intuitive Human Predictions
Model Architecture. As described in Section III, each trajectory $\xi$ is an interaction between two agents. Agents may demonstrate different behaviors during different interactions. Accordingly, we want to learn a low-dimensional space of behaviors $\mathcal{Z}$ that can distinctly represent the different types of behaviors demonstrated by the robot. To this end, we use a special class of autoencoders with finite scalar quantization [31]: instead of a continuous latent space, finite scalar quantization embeds trajectories to a discrete latent space with a fixed number of points. When trained, each of the discrete latent vectors corresponds to a different class of behavior demonstrated by the robot (see Figure 2). Consider a latent space of dimension $d$, where the autoencoder employs a set of $L$ distinct values within the interval $[-1,1]$ to quantize each of the $d$ channels. Each channel of the latent vector can have only one of the $L$ values, and the entire latent space has $L^{d}$ discrete vectors. Mathematically, for a latent vector $z$, each channel $z_{i}$ is linearly scaled to an integer in the range $[0,L-1]$ with the equation:
@@
(see [31] for details). This integer is then re-scaled to a discrete value within the range $[-1,1]$ by applying the formula:
@@
In practice, this scaling and re-scaling maps each channel of the embedding $z$ to a single value in $\{-1,-1+\frac{2}{L-1},\dots,1\}$. For example, if we let $d=3$ and $L=2$, then the discrete latent space will contain $82^{3}=$ 8 vectors. Each channel in this vector can have a value of $-1$ or $1$; therefore, the discrete latent space will be a set $\mathcal{Z}=\{(-1,-1,-1),(-1,-1,1),(-1,1,1),\dots,(1,1,1)\}$.
@@
This solution enables us to take advantage of the quantized latent space. Here, discretization provides a significant bottleneck in the information flow from encoder to decoder (see Figure 2) by limiting the number of latent vectors available for encoding. Consequently, to minimize the loss function, the learned latent space must spread information across the quantized bins: i.e., each latent value should capture a different behavior pattern. Returning to our driving example, in Figure 2, three distinct behaviors of the robot — driving straight, merging left, merging right — are mapped to different latent vectors $z\in\mathcal{Z}$. More generally, quantization of the latent space to discrete values has been shown to avoid posterior collapse — where the latent values fail to capture meaningful representations [30, 31, 35].
@@ V Testing the Encoded High-Level Behaviors
In this section we explore the types of high-level behaviors that are extracted by our approach. Here we do not use actual human data; instead, we consider synthetic interactions, and study whether the underlying patterns our algorithm recovers from these interactions could match with intuitive human interpretations. We consider two multi-agent environments with a human and robot: Highway and Obstacle.
@@
VI User Study
Our ultimate goal is to enable a robot to estimate how humans predict it will behave. In Section V we found that our proposed approach was able to identify the high-level behavior demonstrated by the robot. Now, in this section we validate whether the robot’s extracted behaviors align with those that humans attribute to the robot.
@@ VI User Study
Participants and Procedure. We conducted an online user study with $22$ participants recruited from the Virginia Tech and University of Virginia communities. Every participant completed $10$ trials in each of the environments. In each trial, they observed 4 full interactions, followed by a partial segment of an interaction. After viewing these trajectories, participants predicted the robot’s next $3$ actions, starting from $5$ different initial states, denoted $s_{1}$. Additionally, both Ours and VAE models were used to map the interaction segment to a high-level behavior $z$ within the model’s representation space. The decoder in each model then predicted the actions for the $5$ states using the corresponding $z$.
Hypothesis. Our hypothesis is that:
@@
Ours will more accurately match the predictions made by human participants compared to VAE.
@@
Results. Our results are summarized in Figure 4. The plots in the top-left illustrate the average alignment error across all participants in both environments. We observe that Ours achieves a lower alignment error compared to VAE in both the Highway and Obstacle. A two-tailed paired t-test confirms that these differences in the error are statistically significant ($p<0.01$). This demonstrates that our method is more accurate at estimating how participants predicts the robot will move. This finding supports our hypothesis.
Next, we qualitatively compare the high-level behaviors extracted by Ours with the behaviors attributed to the robot by participants. Specifically, we evaluate whether the intuitive explanations of robot behavior, corresponding to different latent vectors $z$, are consistent with empirical observations. To this end, we plot the robot’s predicted for each $z$ alongside the participants’ predictions, as shown in Figure 4. The top-right plots present the results for Highway, while the bottom row plots display the results for Obstacle.
From the vector fields, we can intuitively infer the robot’s behavior. In the Highway environment, the vector fields indicate that the robot either merges left or merges right, which aligns with the participants’ predictions. Similarly, in the Obstacle environment, the vector fields converge toward specific goals, and participants’ predictions follow the same pattern for each latent vector.
@@
Limitations. While our method provides promising results, it does not perfectly cluster distinct interaction patterns into separate latent vectors. This is evident in the second latent vector $z$ in the Obstacle environment — where our model predicts the robot moving toward the bottom-right goal, participants’ predictions suggest that the robot sometimes moves toward the bottom-left goal. Although the results are encouraging, further work is required to develop more sophisticated second-order Theory of Mind (ToM) models for robots, enabling them to reason more effectively about human mental models.
VII Conclusion
@@
We have presented a theory of mind approach that enables a robot to estimate how humans predict it will behave. Unlike previous research — which often assumes that humans will precisely infer the robot’s policy — we recognized that humans typically rely on course, high-level predictions. For instance: instead of predicting the exact speed and velocity of an autonomous car, the human might reason over whether that car is passing or merging. To extract these high-level predictions directly from interaction data we introduced an autoencoder with a discrete latent space. This latent space served as an information bottleneck (providing course predictions), and its discrete structure yielded different types of behavior (capturing high-level patterns). We tested the feasibility of our method to accurately estimate human predictions through a two-part online user study. Our results suggest that separating the robot’s behavior into high-level clusters leads to more accurate estimates than precise baselines.

### v2

@@ Abstract
Humans interacting with robots often form predictions of what the robot will do next. For instance, based on the recent behavior of an autonomous car, a nearby human driver might predict that the car is going to remain in the same lane. It is important for the robot to understand the human’s prediction for safe and seamless interaction: e.g., if the autonomous car knows the human thinks it is not merging — but the autonomous car actually intends to merge — then the car can adjust its behavior to prevent an accident. Prior works typically assume that humans make precise predictions of robot behavior. However, recent research on human-human prediction suggests the opposite: humans tend to approximate other agents by predicting their high-level behaviors. We apply this finding to develop a second-order theory of mind approach that enables robots to estimate how humans predict they will behave. To extract these high-level predictions directly from data, we embed the recent human and robot trajectories into a discrete latent space. Each element of this latent space captures a different type of behavior (e.g., merging in front of the human, remaining in the same lane) and decodes into a vector field across the state space that is consistent with the underlying behavior type. We hypothesize that our resulting high-level and course predictions of robot behavior will correspond to actual human predictions. We provide initial evidence in support of this hypothesis through proof-of-concept simulations, testing our method’s predictions against those of real users, and experiments on a real-world interactive driving dataset.
@@ I Introduction
This concept of reasoning over another agent’s mental model is known as Theory of Mind (ToM) [5]. Specifically, in this paper we will focus on building a second-order ToM of the human (e.g., "the robot thinks about how human predicts the robot will act"). Existing works often focus only on a first-order ToM: modeling the human’s goal so they can provide better assistance [6], estimating the human’s strategy in order to influence behavior [7, 8], or modeling the human’s intent to improve collaboration between humans and robots [9, 10]. Prior research has also explored building second-order ToM for robots; for instance, [11] models how robot actions affect a human’s behavior and leverage this to develop efficient and influential robot policies. When applied to our example, these works might expect the human to precisely predict each individual action of the autonomous car (see Figure 1, left). However, recent research on human-human interaction suggests that humans do not precisely anticipate the actions of other agents [12]; instead, humans appear to reason over more general and intuitive patterns of behavior. We therefore hypothesize that:
@@
In this paper, we apply our hypothesis to create an algorithmic, data-driven method to estimate how humans predict a robot will behave. Specifically, we develop an autoencoder that inputs recent human and robot behaviors and embeds these trajectories into a discrete latent space (Section IV). This learned latent space (a) inherently forces an information bottleneck that leads to coarse, human-like predictions, and (b) classifies the input trajectories into the discrete, high-level behaviors that were observed in the training dataset. We show, these discrete latent values often correspond to human-friendly patterns: e.g., merging in front of the human, remaining in the same lane, or merging behind the human (Section V). When we decode these discrete latent representations, the robot obtains a vector field of predictions over the environment’s state space. In practice, this enables the robot to estimate how the human thinks it will move right now (in its current state) and also during future timesteps (at other environment states). Our initial user study suggests that the second-order ToM predictions obtained using this vector field better align with actual human predictions than existing baselines (Section VI). Additionally, we validate our model’s ability to capture high-level movement patterns across a large, real-world dataset for driving domains (Section VII).
@@
Testing on Real-World Agent Interactions. We evaluate our proposed method using a real-world interactive driving dataset. The experimental results demonstrate that our approach effectively predicts the robot’s movement by extracting high-level, human-friendly behavior patterns.
@@ II Related Work
Modeling Humans. Existing research in human-robot interaction has explored how robots can model human behavior. These works typically focus on estimating the human’s goal, objective, or intent. For example, in shared autonomy robots infer the human’s goals [13, 14], their intentions [6], even their skills [15] in order to assist humans. Alternatively, in multi-agent interaction robots form high-level representations of humans [7, 16, 17] to model their behavior (e.g., inferring whether human are aggressive or defensive drivers). Overall, these works form first-order models of the human (i.e., they estimate what the human is thinking). But they do not consider how the human reasons over the robot — i.e., how the human expects the robot to behave.
@@ III Problem Statement
Two-Player Game. Formally, the human and robot are playing a two-player game. Let $s_{i}\in\mathcal{S}$ be the state of agent $i$, and let $a_{i}\in\mathcal{A}$ be the action of agent $i$. In what follows, we will use $i=1$ to refer to the robot and $i=2$ to refer to the human. Both agents (i.e., the human and robot) have policies that map the system state to actions. The robot’s policy $\pi_{1}(s)\rightarrow a_{1}$ determines the robot’s actions, and similarly the human’s policy $\pi_{2}(s)\rightarrow a_{2}$ determines the human’s actions. Within our problem setting the robot knows its own policy, but it does not know how the human is modeling the robot’s policy. On the other hand, the human is unaware of the robot’s policy $\pi_{1}$, though they have an understanding of their own predictions regarding the robot’s behavior.
@@ IV Extracting High-Level Behaviors for Intuitive Human Predictions
Model Architecture. As described in Section III, each trajectory $\xi$ is an interaction between two agents. Agents may demonstrate different behaviors during different interactions. Accordingly, we want to learn a low-dimensional space of behaviors $\mathcal{Z}$ that can distinctly represent the different types of behaviors demonstrated by the robot. To this end, we use a special class of autoencoders with finite scalar quantization [31]: instead of a continuous latent space, finite scalar quantization embeds trajectories into a discrete latent space with a fixed number of points. When trained, each of the discrete latent vectors corresponds to a different class of behavior demonstrated by the robot (see Figure 2). Consider a latent space of dimension $d$, where the autoencoder employs a set of $L$ distinct values within the interval $[-1,1]$ to quantize each of the $d$ channels. Each channel of the latent vector can have only one of the $L$ values, and the entire latent space has $L^{d}$ discrete vectors. Mathematically, for a latent vector $z$, each channel $z_{i}$ is linearly scaled to an integer in the range $[0,L-1]$ with the equation [31]:
@@
This integer is then re-scaled to a discrete value within the range $[-1,1]$ by applying the formula:
@@
In practice, the scaling and re-scaling maps each channel of the embedding $z$ to a single value in $\{-1,-1+\frac{2}{L-1},\dots,1\}$. For example, if we let $d=3$ and $L=2$, then the discrete latent space will contain $82^{3}=$ 8 vectors. Each channel in this vector can have a value of $-1$ or $1$; therefore, the discrete latent space will be a set $\mathcal{Z}=\{(-1,-1,-1),(-1,-1,1),(-1,1,1),\dots,(1,1,1)\}$.
@@
This solution enables us to take advantage of the quantized latent space. Importantly, discretization provides a significant bottleneck in the information flow from encoder to decoder (see Figure 2) by limiting the number of latent vectors available for encoding. Consequently, to minimize the loss function, the learned latent space must spread information across the quantized bins: i.e., each latent value should capture a different behavior pattern. Returning to our driving example, in Figure 2, three distinct behaviors of the robot — driving straight, merging left, merging right — are mapped to different latent vectors $z\in\mathcal{Z}$. More generally, quantization of the latent space to discrete values has been shown to avoid posterior collapse — where the latent values fail to capture meaningful representations [30, 31, 35].
@@ V Testing the Encoded High-Level Behaviors
In this section we explore the types of high-level behaviors that are extracted by our approach. Here we do not use actual human data; instead, we start with synthetic interactions, and study whether the underlying patterns our algorithm recovers from these interactions match with intuitive human interpretations. We consider two multi-agent environments with a human and robot: Highway and Obstacle.
@@
VI Comparing to User Predictions
Our ultimate goal is to enable a robot to estimate how humans predict it will behave. In Section V we qualitatively found that our proposed approach was able to identify high-level behaviors demonstrated by the robot. However, so far we have only worked with simulated data. Now in Section VI we put our predictive algorithm to the test with actual humans. We validate whether the robot’s extracted behaviors align with those that real participants attribute to the robot.
@@ VI Comparing to User Predictions
Participants and Procedure. We conducted an online user study with $22$ participants recruited from the Virginia Tech and University of Virginia communities. Every participant completed $10$ trials in each of the environments. In each trial, they observed 4 full interactions, followed by a partial segment of an interaction. After viewing these trajectories, participants predicted the robot’s next $3$ actions, starting from $5$ different initial states, denoted $s_{1}$. Additionally, both Ours and VAE were used to map the interaction segment to a high-level behavior $z$ within the model’s representation space. The decoder in each model then predicted the actions for the $5$ states using the corresponding $z$.
Hypothesis. Our hypothesis H is that:
@@
Ours will more accurately match the predictions made by human participants as compared to VAE.
@@
Results. Our results are summarized in Figure 4. The plots in the top-left illustrate the average alignment error across all participants in both environments. We observe that Ours achieves a lower alignment error compared to VAE in both the Highway and Obstacle. A paired t-test confirms that these differences in the error are statistically significant ($p<0.01$). This demonstrates that our method is more accurate at estimating how participants predicts the robot will move, and supports our hypothesis.
Figure 5: Results of our experiments on the INTERACT dataset [37]. We observe the trajectory of two cars, and predict the movement of one vehicle from the perspective of the other car. We compare the predictions of our high-level method and the precise baseline. (Left) The average alignment error of the predictions across $10$ trials. The left plot shows the results for the Roundabout scenario and the right shows the results for the Intersection scenario. The error value ranges from 0 (indicating parallel prediction) to 2 (indicating opposite prediction). Our method achieves a significantly lower error in both scenarios ($p<0.001$). Next, we visualize the movement patterns identified by the latent vectors of our model. (Top-right) In Roundabout we obtain three prominent movement patterns: entering the roundabout from the left and going up, going around the roundabout, and entering the roundabout from the bottom. (Bottom-right) In Intersection our model identifies three prominent movement patterns: going straight, going left, and going right.
VII Testing on a Real-World Dataset
In Section VI, we validated our approach with data from $22$ participants in custom environments. But just because our approach was able to accurately predict how humans thought a robot would behave in these custom environments does not mean that our approach will work across general, real-world interactions. To more broadly test our approach, in Section VII we leverage an existing dataset that includes driving motions from users around the world. We again explore whether our high-level prediction method is able to infer how humans think other agents will behave.
Dataset. We used the INTERACTION dataset [37] for our analysis. This dataset consists of traffic across a variety of interactive real-world scenarios. Specifically, the dataset provides $4$ second trajectories of cars driving in different road settings collected from $11$ locations. We studied two interactive driving scenarios from this dataset: Roundabout and Intersection. Each trajectory is composed of a sequence of $(x,y)$ positions and velocities for the cars.
Procedure. The INTERACTION dataset provides training and validation sets. We used the training set to learn a model of the agent’s behaviors using our method (Ours) and the baseline (VAE). In each trajectory, we designate one agent as the “robot,” (i.e., the agent whose movement we want to predict), and we designate the other agent as the human. The models observe a $1$ second segment of the trajectory and predict the robot’s next action conditioned on the two agents’ states. We applied the alignment error defined in Section VI to compare the performance of the two methods.
Results. The plot in Figure 5 (left) shows the mean alignment error across $10$ trials. The alignment error for Ours is consistently lower than that of the baseline VAE in both scenarios, Roundabout and Interaction. A paired t-test confirms that these differences in the error are statistically significant ($p<0.001$), indicating that our method more accurately models the robot’s behavior and predicts it’s movements. This finding supports our hypothesis H.
To better understand why our approach was able to generate predictions that better aligned with the actual human’s prediction, we analyzed the high-level patterns extracted by our method. See Figure 5 (right). As a reminder, our goal is to enable the robot to understand how a human thinks it will behave. With the VAE baseline, the robot treats the human as if it were a robot; i.e., the robot thinks the human will make precise, highly accurate predictions. But with Our approach the robot recognizes that human’s reason over higher-level models — instead of precise predictions, the human thinks about how the autonomous car might change lanes or exit the roundabout. Our method autonomously extracts these higher-level predictions in Figure 5. For instance, in the Roundabout scenario the latent vector $z_{1}$ represents the movement of entering the roundabout from the left and heading upwards, while the latent vector $z_{2}$ captures the behavior of navigating around the entire roundabout. Similarly, in the Intersection scenario, different latent vectors correspond to distinct movement patterns: going straight ($z_{1}$), going left ($z_{2}$), or going right ($z_{3}$). These high-level patterns explain why our method is able to more accurately reflect the human predictions across this real-world dataset.
@@ VII Testing on a Real-World Dataset
Limitations and Future Work. While our method yields promising results, it does not perfectly cluster distinct interaction patterns into separate latent vectors. This is evident from the second latent vector $z$ in the Obstacle environment (Figure 4): our model predicts the robot moving toward the bottom-right goal, yet participants’ predictions sometimes indicate movement towards the bottom-left goal. We hypothesize that this limitation could be addressed by incorporating priors or added supervision that captures the types of high-level behaviors humans expect the system to extract.
Moving forward, our future work will incorporate the human’s prediction of the robot’s motion into the robot’s decision making. With our method, the robot now has a data-driven way to express how nearby humans predict that robot will behave. A ToM robot should then reason over this prediction when making its decisions; for instance, if an autonomous car knows that nearby humans expect it to exit the roundabout, and it actually wants to stay on the roundabout, then it should take additional actions to make its intention clear to the humans.
VIII Conclusion
@@
We have presented a theory of mind approach that enables a robot to estimate how humans predict it will behave. Unlike previous research — which often assumes that humans will precisely infer the robot’s policy — we recognized that humans typically rely on course, high-level predictions. For instance: instead of predicting the exact speed and velocity of an autonomous car, the human might reason over whether that car is passing or merging. To extract these high-level predictions directly from interaction data we introduced an autoencoder with a discrete latent space. This latent space served as an information bottleneck (providing course predictions), and its discrete structure yielded different types of behavior (capturing high-level patterns). We tested the feasibility of our method to accurately estimate human predictions through i) proof-of-concept simulations, ii) a user study with real human predictions, and iii) tests across a large, real-world interactive driving dataset. Our results suggest that separating the robot’s behavior into high-level clusters leads to more accurate estimates than precise baselines.